# Bourbaki degree of pairs of projective surfaces

### 31 Mar 2026 v3

@@ 2.3 Free resolutions and the Bourbaki scheme
and in particular $\sigma$ is a nearly free sequence.
@@
If $\sigma=(f,g)$ is a nearly free normal sequence, then ${\mathcal{T}}_{\sigma}$ is not locally free.
@@
which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$, hence ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable.
@@
which yields
$m_{0}-m(\sigma)-\frac{d^{2}}{4}=0,$
so that $c_{2}(E)=0$. But a $\mu$-semistable reflexive sheaf with $c_{1}(E)=c_{2}(E)=0$ must be $E\simeq\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 2}$ ([undefn, Lemma 9.7]), implying $\sigma$ is free, contradicting the hypothesis. ∎
@@
If a normal sequence $\sigma=(f,g)$ is a $3$-syzygy, then $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$. Moreover, a sequence $\sigma$ is $3$-syzygy if and only if $B_{\nu}$ is a complete intersection, for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
@@
$0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\lambda}{\mathcal{T}}_{\sigma}\rightarrow 0,$
so we split the resolution into two short exact sequences:
$F_{2}\hookrightarrow F_{1}\twoheadrightarrow S\text{ and }S\hookrightarrow F_{0}^{\prime}\twoheadrightarrow{\mathcal{T}}_{\sigma},$
and focus on the second one. The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [undefn, Proposition 1.1]. Furthermore, since $\operatorname{{rk}}(F_{0}^{\prime})=3$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
@@
For the equivalence stated above, if we start with a sequence $\sigma$ which is $3$-syzygy and then apply Lemma 12, $(b)$, we obtain a resolution for $\mathcal{I}_{B}$ which is of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\rightarrow\mathcal{I}_{B}\rightarrow 0,$
since $\operatorname{{rk}}(F_{0})=2$, thus concluding $B$ must be a complete intersection scheme. On the other hand, if $B$ is a complete intersection, then there is a resolution for $\mathcal{I}_{B}$ of the form
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-k)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-l)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-d)\xrightarrow{r,s}\mathcal{I}_{B}\rightarrow 0,$
induced by the two equations $r,s$ defining $B=V(r,s)\subset\mathbb{P}^{3}$. Thus, applying Lemma 12, $(a)$, we can lift the resolution above for a free resolution for ${\mathcal{T}}_{\sigma}$ of the form
$0\rightarrow F_{1}\rightarrow F_{0}^{\prime}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0$
so that $\operatorname{{rk}}(F_{0}^{\prime})=3$, thus $\operatorname{{rk}}(F_{0})=2$ and $\sigma$ is $3$-syzygy. ∎
@@
where the first follows from 20 since $\sigma$ is nearly free iff $B_{\nu}$ is a line for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and every line is a complete intersection of two planes. The converses do not hold, as we explore in the next examples: there are $3$-syzygy pencils of cubics which are not nearly free (Example 15), pencils of cubics which satisfy $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but are not $3$-syzygy (Example 22) and pencils of cubics with $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (Example 23).
We also construct two pencils of cubics with the same discrete invariants $(m(\sigma),\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma))$ and the same Chern classes, which are distinguished by their homological behavior: one is $3$-syzygy and the other satisfies $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$ (see Example 22 and Example 24).
Example 22 (pencil of cubics which is not $3$-syzygy and $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$).
Considering the following pencil of cubics:
@@
The scheme $(\Xi_{\sigma})_{1}$ has three irreducible components, given by two lines and a plane quadric curve:
$L_{1}$ $=V(x_{0},x_{3})\doteq V(\mathfrak{p}_{1})$
$L_{2}$ $=V(x_{2},x_{3})=V(\mathfrak{p}_{2})$
$Q$ $=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})=V(\mathfrak{p}_{3}),$
Now, we proceed to computing the length of the cokernel module $Q$ over each prime ideal $\mathfrak{p}_{i}$, in order to show that $m(\sigma)=5$ using the associativity formula.
Starting with the prime ideal $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that the element $3x_{2}^{2}\in R_{\mathfrak{p}_{1}}$ is invertible, and hence the sum $v=3x_{2}^{2}+x_{1}x_{3}$ is also invertible, as $x_{1}x_{3}\in\mathfrak{p}_{1}$. Let $u=v^{-1}$, so we perform two elementary operations on the matrix $\nabla\sigma$ over $\mathfrak{p}_{1}$:
$(\nabla\sigma)_{\mathfrak{p}_{1}}$ $\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&x_{2}x_{3}&v&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$
$\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&ux_{2}x_{3}&1&u(x_{1}x_{2}+3x_{3}^{2})\end{pmatrix}$.
The first transform is to add the scaled second row (by $ux_{0}^{2}$) to the first, and then we scale the second row by $u$. Then, the third basis vector goes via $\nabla\sigma$ to $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, and therefore we may compute the cokernel from the quotient
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(2x_{0}x_{2}+x_{1}x_{3},x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3},x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2}))}\simeq\frac{R_{\mathfrak{p}_{1}}}{I},$
where $I=(f_{1},f_{2},f_{3})$ are the generators above. We may factor $f_{2}=x_{0}x_{3}(1-ux_{0}x_{2})$, so that since $(1-ux_{0}x_{2})$ is a unit in the local ring, we conclude $x_{0}x_{3}\in I$. From this, multiplying $x_{3}f_{1}$ we conclude that $x_{1}x_{3}^{2}\in I$, and since $x_{1}$ is also a unit, we obtain $x_{3}^{2}\in I$, and analogously taking $x_{0}f_{1}$ we obtain $x_{0}^{2}x_{2}\in I$, and since $x_{2}$ is a unit we get $x_{0}^{2}\in I$. Thus
$(x_{0}^{2},x_{0}x_{3},x_{3}^{2})\subset I.$
Now, from this, looking at $f_{3}$:
$f_{3}=x_{0}x_{1}+3x_{3}^{2}-x_{0}^{2}u(x_{1}x_{2}+3x_{3}^{2})\in I\Rightarrow x_{0}x_{1}\in I\Rightarrow x_{0}\in I,$
and looking at $f_{1}$, we obtain
$f_{1}=2x_{0}x_{2}+x_{1}x_{3}\in I\Rightarrow x_{1}x_{3}\in I\Rightarrow x_{3}\in I,$
so that $I=(x_{0},x_{3})$ and thus $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\kappa$, with $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
For the case $\mathfrak{p}_{2}=(x_{2},x_{3})$, let $u=(x_{0}^{2})^{-1}$, $q=(3x_{2}^{2}+x_{1}x_{3})$ and consider the elementary operations given by $R_{2}^{\prime}=R_{2}-uqR_{1}$ and then $R_{1}^{\prime}=uR_{1}$:
$(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\sim\begin{pmatrix}u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&u(x_{0}x_{1}+3x_{3}^{2})\\ -uq(2x_{0}x_{2}+x_{1}x_{3})&x_{2}x_{3}-uqx_{0}x_{3}&0&x_{1}x_{2}+3x_{3}^{2}-uq(3x_{3}^{2}+x_{0}x_{1})\end{pmatrix}$.
The Jacobian matrix sends the third basis vector to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, and thus we may compute:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I}$
where
$f_{1}$ $=(3x_{2}^{2}+x_{1}x_{3})(2x_{0}x_{2}+x_{1}x_{3})$
$f_{2}$ $=x_{3}(x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3})$
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3ux_{3}^{2}(3x_{2}^{2}+x_{1}x_{3})-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})$
First, we consider
$f_{2}-\frac{x_{3}}{x_{1}}f_{3}=\frac{3x_{3}^{3}}{x_{1}}(1-u(3x_{2}^{2}+x_{1}x_{3}))\in I,$
and thus $x_{3}^{3}\in I$. Then
$x_{3}f_{2}=x_{3}^{2}x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3}^{3}\in I\Rightarrow x_{3}^{2}x_{2}\in I,$
and moreover
$x_{3}f_{1}$ $=x_{2}x_{3}(6x_{0}x_{2}^{2}+3x_{1}x_{2}x_{3}+2x_{0}x_{1}x_{3})+x_{1}^{2}x_{3}^{3}\in I$
$\Rightarrow 6x_{0}x_{2}^{3}x_{3}+x_{3}^{2}x_{2}(3x_{1}x_{2}+2x_{0}x_{1})\in I$
$\Rightarrow 6x_{0}x_{2}^{3}x_{3}\in I$
$\Rightarrow x_{2}^{3}x_{3}\in I.$
On the other hand,
$x_{2}x_{3}f_{3}$ $=-3ux_{3}^{3}q+3x_{2}x_{3}^{3}+x_{2}x_{3}(x_{1}x_{2}-ux_{0}x_{1}q)\in I$
$\Rightarrow x_{1}x_{2}^{2}x_{3}-3ux_{0}x_{1}x_{2}^{3}x_{3}-ux_{0}x_{1}^{2}x_{2}x_{3}^{2}\in I$
$\Rightarrow x_{1}x_{2}^{2}x_{3}\in I\Rightarrow x_{2}^{2}x_{3}\in I.$
Now, consider
$x_{3}f_{3}$ $=x_{1}x_{2}x_{3}+3x_{3}^{3}-uqx_{3}^{3}-uqx_{0}x_{1}x_{3}\in I$
$\Rightarrow x_{3}x_{2}(x_{2}-3ux_{0}x_{2}^{2}-ux_{0}x_{1}x_{3})\in I$
$\Rightarrow x_{1}x_{3}(x_{2}-ux_{0}x_{1}x_{3})\doteq A\in I,$
and
$x_{3}A\in I\Rightarrow x_{1}x_{3}^{2}x_{2}-ux_{0}x_{1}^{2}x_{3}^{2}\in I\Rightarrow x_{3}^{2}\in I.$
This simplifies
$f_{2}=x_{3}(x_{2}-uqx_{0})\in I\Rightarrow x_{2}x_{3}(1-3ux_{2})\in I\Rightarrow x_{2}x_{3}\in I.$
Considering
$f_{1}$ $=6x_{0}x_{2}^{3}+x_{2}x_{3}(3x_{1}x_{2}+2x_{0}x_{1})+x_{1}^{2}x_{3}^{2}\in I$
$\Rightarrow x_{2}^{3}\in I,$
and then
$x_{2}f_{3}$ $=x_{1}x_{2}^{2}+x_{2}(3x_{3}^{2}-3ux_{3}^{2}q)-uqx_{0}x_{1}x_{2}\in I$
$\Rightarrow x_{1}x_{2}^{2}-3ux_{0}x_{1}x_{2}^{3}-ux_{0}x_{1}^{2}x_{2}x_{3}\in I$
$\Rightarrow x_{1}x_{2}^{2}\in I\Rightarrow x_{2}^{2}\in I,$
simplifying
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3uqx_{3}^{2}-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})\in I$
$\Rightarrow x_{1}(x_{2}-ux_{0}x_{1}^{2}x_{3})\in I\Rightarrow x_{2}-ux_{0}x_{1}x_{3}\in I,$
so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa\langle 1,x_{2}\rangle$, meaning $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=2$, since the equations $f_{1},f_{2}$ are trivialized after $x_{3}^{2},x_{2}^{2},x_{2}x_{3}\in I$.
Let us consider the case $\mathfrak{p}_{3}=(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})\doteq(g_{1},g_{2})$. One may be tempted to choose $u=(2x_{0}x_{2}+x_{1}x_{3})^{-1}$ and do the elementary operations $R_{1}^{\prime}=uR_{1}$, but
$2x_{0}x_{2}+x_{1}x_{3}=2x_{0}g_{1}+g_{2}.$
Take $u\doteq(x_{0}^{2})^{-1}\in R_{\mathfrak{p}_{3}}$. If $x_{0}^{2}\in\mathfrak{p}_{3}$, then:
$x_{0}\in\mathfrak{p}_{3}\Rightarrow x_{2}\in\mathfrak{p}_{3}\Rightarrow x_{1}x_{3}\in\mathfrak{p}_{3}$
and $\text{height}(\mathfrak{p}_{3})\geq 3$ would give a contradiction. Thus, we may perform the same elementary operations as we did for $\mathfrak{p}_{2}$, and thus
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\frac{R_{\mathfrak{p}_{3}}}{I}$
with generators $I=(f_{1},f_{2},f_{3})$ given as before. By the same reasoning, $(3x_{2}^{2}+x_{1}x_{3})\not\in\mathfrak{p}_{3}$, and thus
$f_{1}\in I\Rightarrow 2x_{0}x_{2}+x_{1}x_{3}\in I.$
From the previous description of the element above, it suffices to show $g_{i}\in I$ for some $i\in\{1,2\}$ to conclude $I=\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$. Let $A\doteq(g_{2}+x_{2}^{2})$, so that $u=A^{-1}$. We can rewrite $f_{3}$ as:
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-u(g_{2}+x_{2}^{2})(3x_{3}^{2}+x_{0}x_{1})$
$=3x_{3}^{2}(1-u(g_{2}+x_{2}^{2}))+x_{1}(x_{2}-ux_{0}(g_{2}+x_{2}^{2}))$
$=3x_{3}^{2}(1-1)+x_{1}(x_{2}-ux_{0})$
$=-x_{1}g_{1}\in I,$
and since $x_{1}\notin\mathfrak{p}_{4}$, $g_{1}\in I$ and $I=\mathfrak{p}_{4}R_{\mathfrak{p}_{4}}$, concluding $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=1$. From the associativity formula, we obtain
$m(\sigma)=1\cdot 1+2\cdot 1+1\cdot 2=5.$
Since $m(\sigma)\leq 6$, we conclude using the bounds from the next section that $e\geq 3$ The minimal free resolution of ${\mathcal{T}}_{\sigma}$ given using Macaulay2 is
@@
considered in [undefj, Theorem 8.1]. From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle. Therefore, we obtain $\operatorname{Bour}(\sigma)=2$, and a free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
@@
So $e=3$ and $m(\sigma)=7$. The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$. Localizing the cokernel at the first two lines gives length one, so over last line should have length $5$.
Example 24 (pencils of cubics with the same discrete invariants and different homological behavior).
Considering the following pencil of cubics:
@@
Here, the unique irreducible component of $(\Xi_{\sigma})_{1}$ is the line $V(x_{0},x_{3})$.
Moreover, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
@@
This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as Example 15, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
In Section 4 we characterize all nearly free pencils of cubics $(d_{f}=d_{g}=2)$ and all nearly free sequences with $d_{f}=1,d_{g}=2$.

### 27 Apr 2026 v4

@@ 2.3 Free resolutions and Bourbaki schemes
and in particular $\sigma$ is a nearly-free sequence.
@@
If $\sigma=(f,g)$ is a nearly-free normal sequence, then ${\mathcal{T}}_{\sigma}$ is not locally free.
@@
which implies that $d$ must be even and that $e=d/2+1$, giving $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$, hence ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable.
@@
which yields $m_{0}-m(\sigma)-\frac{d^{2}}{4}=0$, so that $c_{2}(E)=0$. But a $\mu$-semistable reflexive sheaf with $c_{1}(E)=c_{2}(E)=0$ must be $E\simeq\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 2}$ (by [14, Lemma 9.7]), implying $\sigma$ is free, contradicting the hypothesis. ∎
@@
If a normal sequence $\sigma=(f,g)$ is a $3$-syzygy, then $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$. Moreover, a sequence $\sigma$ is $3$-syzygy if and only if $B_{\nu}$ is a complete intersection, for $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))\setminus\{0\}$, $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$.
@@
$0\rightarrow F_{2}\rightarrow F_{1}\rightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\xrightarrow{\lambda}{\mathcal{T}}_{\sigma}\rightarrow 0$
with $\operatorname{{rk}}(F_{0})=2$. Split the resolution into two short exact sequences:
$F_{2}\hookrightarrow F_{1}\twoheadrightarrow S\text{ and }S\hookrightarrow F_{0}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-e)\twoheadrightarrow{\mathcal{T}}_{\sigma},$
and focus on the second one. The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [14, Proposition 1.1]. Furthermore, since $\operatorname{{rk}}(F_{0})=2$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
@@
With this and 13, we conclude that $\sigma$ is $3$-syzygy if and only if $B=B_{\nu}$ admits a resolution of the form:
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-k)\to\mathcal{O}_{\mathbb{P}^{3}}(-a)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-b)\to\mathcal{I}_{B}\to 0,$
which is equivalent to the fact that $B\subset\mathbb{P}^{3}$ is a complete intersection. ∎
@@
from 17 and 21. The converses do not hold, as we explore in the examples: there are $3$-syzygy pencils of cubics which are not nearly free (Example 16), and the next example is a pencil of cubics satisfying $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but not $3$-syzygy.
Example 23 ($\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$, not $3$-syzygy).
Consider the following pencil of cubics:
@@
The scheme $(\Xi_{\sigma})_{1}$ has three irreducible components, given by two lines $V(x_{0},x_{3}),V(x_{2},x_{3})$ and a plane quadric curve $V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})$. Using the associativity formula, one may compute $m(\sigma)=5$.
The minimal free resolution of ${\mathcal{T}}_{\sigma}$ given using Macaulay2 is
@@
considered in [11, Theorem 8.1]. From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle. Therefore, we obtain $\operatorname{Bour}(\sigma)=2$, and a free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
@@
So $e=3$ and $m(\sigma)=7$. The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$.
For normal sequences $\sigma=(f,g)$, the triples $(\operatorname{indeg}({\mathcal{T}}_{\sigma},\operatorname{Bour}(\sigma),m(\sigma))$ are not enough ensure the homological behavior of the sheaf ${\mathcal{T}}_{\sigma}$. The resolutions for the associated Bourbaki schemes also fix the arithmetic genus $p_{a}(B)$, which describes the third Chern class $c_{3}({\mathcal{T}}_{\sigma})$ of the sheaf. Even if these are fixed, there are curves with fixed degree and genus with different minimal free resolutions. The next example has the same triple $(\operatorname{indeg}({\mathcal{T}}_{\sigma},\operatorname{Bour}(\sigma),m(\sigma))$ as in the $3$-syzygy Example 16 seen before, but this one has $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=2$.
Example 25.
@@
Here, the unique irreducible component of $(\Xi_{\sigma})_{1}$ is the line $V(x_{0},x_{3})$. Using Macaulay2, we obtain a free resolution for ${\mathcal{T}}_{\sigma}$ of the form:
@@
In Section 4 we characterize all nearly-free pencils of cubics $(d_{f}=d_{g}=2)$ and all nearly-free sequences with $d_{f}=1,d_{g}=2$.